OpenStudy (autum0825):

What is the value of x/y when x=9/4 and y=3/5

OpenStudy (triciaal):

you need to be here to get help. Do not just post a question and leave. According to the computer you are offline and the question posted 4 mins ago.

OpenStudy (s4sensitiveandshy):

divide x by y \[\frac{x}{y}\] x is 9/4 , replace x with 9/4 y is 3/5

OpenStudy (s4sensitiveandshy):

and then simplify \[\frac{ \frac{ a }{b } }{ \frac{ c }{ d } } = \frac{a}{b} \times \frac{d}{c}\] multiply numerator by the reciprocal of the denominator